How many electrons pass through an ammeter if 0.0200 amps of current flows for 10.0 seconds?



1
Expert's answer
2021-08-02T08:41:25-0400

The number of electrons that passes through an ammeter 

"N=\frac{\Delta q}{e}=\frac{I\Delta t}{e}""N=\frac{(0.0200\:\rm A)*(10.0\:\rm s)}{1.60\times 10^{-19}\:\rm C}=1.25\times 10^{18}"